if x^2 + y^2 =86 and xy = -16 find the value of (x-y)^2
1
Expert's answer
2013-02-27T05:25:39-0500
Let's transform required expression so that it can be represented in terms of known expressions:(x-y)^2=x^2-2*xy+y^2=(x^2+y^2) +(-2xy)=Now we can substitute given values to obtain the final result:=86+(-2)*(-16)=86+32=118